# Photography Perspective Composition: Towards Aesthetic Perspective Recommendation

### v2

@@ Abstract
Traditional photography composition approaches are dominated by 2D cropping-based methods. However, these methods fall short when scenes contain poorly arranged subjects. Professional photographers often employ perspective adjustment as a form of 3D recomposition, modifying the projected 2D relationships between subjects while maintaining their actual spatial positions to achieve better compositional balance. Inspired by this artistic practice, we propose photography perspective composition (PPC), extending beyond traditional cropping-based methods. However, implementing the PPC faces significant challenges: the scarcity of perspective transformation datasets and undefined assessment criteria for perspective quality. To address these challenges, we present three key contributions: (1) An automated framework for building PPC datasets through expert photographs. (2) A video generation approach that demonstrates the transformation process from suboptimal to optimal perspectives. (3) A perspective quality assessment (PQA) model constructed based on human performance. Our approach is concise and requires no additional prompt instructions or camera trajectories, helping and guiding ordinary users to enhance their composition skills.
@@ 1 Introduction
Professional photography demands expertise in multiple aspects, with photographic composition being one of the most crucial. Photographic composition refers to the arrangement of visual elements according to aesthetic principles. It requires photographers to harmoniously integrate multiple elements like people, urban, and natural features. While master photographers, such as those in Magnum Photos, demand professional knowledge and extensive training, making quality photography expensive and challenging for ordinary people. This raises the question: Can we help ordinary people achieve professional-level composition?
Traditional photography composition approaches are primarily based on cropping. While numerous approaches have been developed for image cropping, including saliency-based methods tu2020image , learning-based techniques deng2018aesthetic ; guo2018automatic ; hong2021composing ; li2020composing ; lu2019listwise ; mai2016composition ; wei2018good ; zeng2019reliable ; zeng2020grid , and reinforcement learning strategies li2018a2 . However, crop-based methods are inherently limited, as they only allow for 2D recomposition within the image plane. As shown in Fig. 1a, traditional crop-based methods primarily focus on learning a crop template. However, when the scene itself is chaotic and not an optimal perspective, cropping alone rarely produces satisfactory results.
@@
In real-world scenarios, photographers address these limitations of 2D cropping by actively adjusting their perspective and positions to achieve optimal spatial relationships between subjects. Through perspective adjustments, photographers can create sophisticated compositions by systematically arranging subjects within the frame, manipulating spatial relationships to create dynamic and engaging images. Fig. 1b illustrates how perspective transformation can address compositional challenges.
@@ 2.1 Photography Composition
Prior photography composition methods primarily rely on image cropping. There are three types. (1) free-form cropping. Numerous techniques have been explored to tackle this problem, including saliency maps tu2020image , learning-based methods deng2018aesthetic ; guo2018automatic ; hong2021composing ; kao2017automatic ; li2020composing ; lu2019listwise ; mai2016composition ; tu2020image ; wei2018good ; zeng2019reliable ; zeng2020grid , and reinforcement learning li2018a2 . (2) Subject-aware image cropping hong2023learning ; yang2023focusing , where an additional subject mask is provided to indicate the subject of interest. (3) Ratio-aware cropping christensen2021experience , where the crops are expected to adhere to a specified aspect ratio. Several datasets have been established, including GAIC gaic , CPC cpc , FCDB fcdb , and SACD sacd . Recent advances in diffusion models have further expanded the possibilities, with works leveraging Stable Diffusion stablediffusion for synthetic data generation fakeitmakeit ; stablerep and developments in outpainting hong2023learning ; ganextend ; verylongscenery . However, cropping methods share a fundamental limitation as they operate only in 2D space, making them insufficient when the spatial arrangement of the scene is suboptimal.
@@ 2.2 Evaluation with Human Performance
Evaluation models are essential for aligning generative models with human preferences. Traditional metrics like FID heusel2017gans and CLIP scores radford2021learning have been widely used huang2023t2i ; huang2024vbench ; liu2024evalcrafter . To improve evaluation accuracy, recent works kirstain2023pick ; liang2024rich ; wu2023humanv2 ; xu2024imagereward ; zhang2024learning have evolved from simple metrics to learning-based approaches that leverage human preference datasets to train CLIP-based models. Recently, researchers have begun exploring vision-language models (VLMs) achiam2023gpt ; wang2024qwen2 as a more powerful framework for reward modeling. These VLM-based approaches have shown success in both evaluation he2024videoscore ; wu2023humanv2 ; wu2023human and optimization lee2023aligning ; prabhudesai2024video ; wallace2024diffusion ; xu2024imagereward , utilizing methods like point-wise regression he2024videoscore ; xu2024visionreward , pair-wise comparison with Bradley-Terry loss bradley1952rank , and instruction tuning li2023generative ; lin2024criticbench ; wang2024lift to leverage reasoning capabilities for VLMs. However, VLMs need substantial data for effective training, yet expert compositional data is scarce and costly to obtain, making it challenging to train VLMs to capture sophisticated aesthetic principles using only limited expert annotations.
@@ 3.2 Automated Construction of PPC Dataset
[Intuition.] Currently, no dedicated dataset exists for PPC setting. To promote this area, we propose a novel approach for constructing PPC dataset in an automated way. The main challenge lies in collecting real-world camera movement sequences that transition from suboptimal to optimal perspective. The closest existing data comes from photographers sharing point-of-view (POV) recordings of their shooting process on streaming platforms. However, these videos are typically captured from secondary angles using GoPro, which differs from the main camera perspective. We observe that expertly composed photographs are readily available unsplash , and recent advances in single-image scene reconstruction have shown remarkable results yu2024viewcrafter . This inspired us to explore an alternative approach: generating camera movement sequences that transition from optimal to suboptimal perspective based on existing well-composed photographs. By reversing these sequences, we can obtain the desired data.
@@
[Detail.] (1) Data Source. We select multiple professional photography datasets, including datasets used in existing composition studies such as GAIC zeng2020grid , SACD yang2023focusing , FLMS flms , and FCDB fcdb . Furthermore, to expand our data volume, we incorporated Unsplash unsplash2023 , currently the largest open-source professional photography dataset. (2) Perspective Transformation Generation. As shown in Fig. 2, we adopt a 3D reconstruction approach. Our 3D reconstruction methodology mainly builds upon the ViewCrafter yu2024viewcrafter . The inputs consist of a well-composed image and a specified camera motion trajectory. Note that this trajectory can be random (refer to [Discussion]). By following this trajectory, we can generate a video sequence transitioning from the optimal to suboptimal perspective. Then, by reversing this video sequence, we obtain our desired training data. (3) Data Filtering. Given the limited performance of current reconstruction models, the generated video data needs to filter out artifacts including distortion, fixedness, and blur effects, as depicted in Fig. 2. However, manual filtering for such a large dataset is impractical. Our tests showed that a single person can only filter about 3K videos per day, making it difficult to process large-scale samples. With the rapid advancement of vision language models (VLMs) in scene understanding and automated evaluation liu2025improving ; prabhudesai2024video ; wu2023human , we develop a perspective quality assessment (PQA) model to filter the generated data. For specific details about the PQA construction, please refer to Sec. 3.4. Our PQA evaluates generated data across three dimensions: visual quality (VQ), motion quality (MQ), and composition aesthetic (CA). These individual scores are aggregated into a final score, and samples exceeding a threshold are selected as our training data. Further details can be found in Appendix D.
@@
[Discussion.] How to handle cases where the initial perspective is suboptimal? In our pipeline, we initially treat the original image as the optimal perspective. This assumption is later refined through human preference learning during the PQA filtering (Sec. 3.4) and RLHF (Sec. 3.3) stages, acknowledging that the original perspective may not always be optimal.
@@ 3.3 Photography Perspective Composition (PPC)
[Intuition.] Previous image composition works primarily focused on cropping approaches. In contrast, we propose a video-based approach: given a suboptimal perspective, we generate a camera movement sequence that gradually transitions to an optimal perspective of the scene. This video-based approach is motivated by two key observations: First, image composition represents a partial order relationship where quality assessment often relies on comparing different views of the same scene rather than isolated evaluation. Second, the transition process naturally demonstrates compositional improvements through before-and-after comparisons, making it particularly effective for both visual demonstration and educational purposes.
[Detail.] (1) Base Pipeline. As shown in Fig. 3, our pipeline takes a suboptimal perspective as input and generates a transformation video from the suboptimal to optimal perspective. This process can be modeled as an image-to-video (I2V) task. I2V has seen remarkable progress, with both commercial solutions like OpenAI’s Sora openaisora2024 , Runway Gen-3 runway2024gen3 , and Pika 1.5 pika2024pika , and open-source models like Hunyuan kong2024hunyuanvideo , CogVideo yang2024cogvideox , and WAN wang2025wan . Our pipeline leverages these open-source models to generate perspective transformation videos. We utilize the last frame of the video as our final optimal perspective and design a method to guide human actions. First, we draw a guidance box (the red bbox in Fig. 3) on the optimal perspective. Then, based on this box, along with the initial and final perspectives, we transform this box onto the original image using feature matching, creating a distorted box. As the user moves, this box gradually changes shape, approaching a rectangle when reaching the true optimal perspective. To simplify the process and accelerate computation, we only use traditional homography transformation dubrofsky2009homography .
@@
(2) RLHF for Quality Enhancement. We observed that some generated perspective transformation videos, while deviating from the GT direction, still maintain high aesthetic quality. This suggests that strict GT adherence may be suboptimal. Therefore, we propose incorporating direct preference optimization (DPO) to align the model with human preferences. This approach encourages the exploration of aesthetically pleasing trajectories that may differ from GT, avoiding the limitation where GT-based optimization could discourage potentially superior compositional alternatives.
@@ 3.4 Perspective Quality Assessment (PQA) Model
[Detail.] (1) Dateset Setting. Stage ①: Unpaired Videos. This stage focuses on distinguishing video quality levels. We collected approximately 5K perspective transformation videos generated by 3D reconstruction models, with expert annotators identifying roughly 1.5K high-quality and 3.5K low-quality samples. To expand the dataset, we randomly paired each high-quality video with 10 low-quality ones, creating a 15K unpaired dataset. Stage ②: Paired Videos. This stage focuses on composition aesthetic recognition through paired comparison learning. For each suboptimal perspective, we generate three video clips using separately trained CogVideoX 1.5, WAN 2.1, and the original GT data. These clips are paired with each other, where "paired" indicates videos sharing identical input views. Expert annotators evaluate these pairs across three dimensions: visual quality (VQ), motion quality (MQ), and composition aesthetic (CA). For each dimension, annotators choose between options (A wins/Ties/B wins). Notably, the CA metric assesses the compositional improvement throughout the video transformation rather than static frame quality.
Figure 4: Quality and human performance results for PPC.
Detailed annotation guidelines are provided in Appendix A.
@@
(2) Model Setting. Our model primarily follows the architecture of VideoAlign liu2025improving . We employ Qwen 2-VL as our base model, utilizing the Bradley-Terry model with ties loss (BTT) rao1967ties , which extends the traditional Bradley-Terry frameworkbradley1952rank . To better handle multi-dimensional evaluation, we separate special tokens for context-agnostic (visual quality, motion quality) and composition-aware (composition aesthetic) attributes, leveraging the causal attention mechanism to achieve effective feature decoupling. The model predicts rewards for each dimension through a shared linear projection head applied to the corresponding token representations from the final layer. For detailed information, please refer to Appendix B.
@@ 4.1 Investigation for Photography Perspective Composition (PPC)
Metric Design. Since there is no prior work on photography composition using perspective transformation, we need to define metrics. We evaluate the generated videos from three parts: perspective accuracy, video quality, and human performance score. Please refer to Appendix E for details.
@@
Main Results. [Quantitative Results] As shown in Tab. 1, we demonstrate the performance of three I2V models in generating PPC videos. [Qualitative Results] We demonstrate the versatility of our approach across three representative scenarios. The first scenario addresses single subjects (e.g., human figures and animals), as shown in Fig. 5, where our PPC model enhances compositional harmony by seamlessly integrating subjects with their surroundings. In the second scenario (Fig. 7), we tackle multi-subject scenes, demonstrating how PPC achieves balanced spatial arrangements to elevate overall visual aesthetics. The third scenario explores landscape photography (Fig. 8), addressing two common challenges faced by novice photographers: balance and horizontal alignment. Our model effectively optimizes these scenes, particularly enhancing symmetrical compositions.
@@
Beyond these scenarios, we discovered the applicability of PPC to UAV photography. As illustrated in Fig. 9, our model successfully identifies optimal views from drone-like perspectives, generating camera movements that adhere to compositional principles while maintaining aesthetic appeal.
The Consistency of PPC. Fig. 6b demonstrates our model’s consistency in PPC. When presented with different suboptimal views of the same scene, our model generates consistent optimal perspectives, maintaining coherence across different inputs.

### v3

@@ Abstract
Traditional photography composition approaches are dominated by 2D cropping-based methods. However, these methods fall short when scenes contain poorly arranged subjects. Professional photographers often employ perspective adjustment as a form of 3D recomposition, modifying the projected 2D relationships between subjects while maintaining their actual spatial positions to achieve better compositional balance. Inspired by this artistic practice, we propose photography perspective composition (PPC), extending beyond traditional cropping-based methods. However, implementing the PPC faces significant challenges: the scarcity of perspective transformation datasets and undefined assessment criteria for perspective quality. To address these challenges, we present three key contributions: (1) An automated framework for building PPC datasets through expert photographs. (2) A video generation approach that demonstrates the transformation process from less favorable to aesthetically enhanced perspectives. (3) A perspective quality assessment (PQA) model constructed based on human performance. Our approach is concise and requires no additional prompt instructions or camera trajectories, helping and guiding ordinary users to enhance their composition skills.
@@ 1 Introduction
Professional photography demands expertise in multiple aspects, with photographic composition being one of the most crucial. Photographic composition refers to the arrangement of visual elements according to aesthetic principles. It requires photographers to harmoniously integrate multiple elements like people, urban, and natural features. Master photographers, such as those in Magnum Photos, require professional knowledge and extensive training, making quality photography expensive and challenging for ordinary people. This raises the question: Can we help ordinary people achieve professional-level composition?
Traditional photography composition approaches are primarily based on cropping. Numerous approaches have been developed for image cropping, including saliency-based methods tu2020image , learning-based techniques deng2018aesthetic ; guo2018automatic ; hong2021composing ; li2020composing ; lu2019listwise ; mai2016composition ; wei2018good ; zeng2019reliable ; zeng2020grid , and reinforcement learning strategies li2018a2 . However, crop-based methods are inherently limited, as they only allow for 2D recomposition within the image plane. As shown in Fig. 1a, traditional crop-based methods primarily focus on learning a crop template. However, when the scene itself is chaotic and lacks good compositional structure, cropping alone rarely produces satisfactory results.
@@
In real-world scenarios, photographers address these limitations of 2D cropping by actively adjusting their perspective and positions to achieve improved spatial relationships between subjects. Through perspective adjustments, photographers can create sophisticated compositions by systematically arranging subjects within the frame, manipulating spatial relationships to create dynamic and engaging images. Fig. 1b illustrates how perspective transformation can address compositional challenges.
@@ 2.1 Photography Composition
Prior photography composition methods primarily rely on image cropping. There are three types. (1) free-form cropping. Numerous techniques have been explored to tackle this problem, including saliency maps tu2020image , learning-based methods deng2018aesthetic ; guo2018automatic ; hong2021composing ; li2020composing ; lu2019listwise ; mai2016composition ; tu2020image ; wei2018good ; zeng2019reliable ; zeng2020grid , and reinforcement learning li2018a2 . (2) Subject-aware image cropping hong2023learning ; yang2023focusing , where an additional subject mask is provided to indicate the subject of interest. (3) Ratio-aware cropping christensen2021experience , where the crops are expected to adhere to a specified aspect ratio. Several datasets have been established, including GAIC gaic , CPC cpc , FCDB fcdb , and SACD sacd . Recent advances in diffusion models stablediffusion ; beta2024zheng have further expanded the possibilities, with works leveraging Stable Diffusion for synthetic data generation fakeitmakeit ; stablerep and developments in outpainting hong2023learning ; ganextend ; verylongscenery . However, cropping methods share a fundamental limitation as they operate only in 2D space, making them insufficient when the spatial arrangement of the scene is less favorable.
@@ 2.2 Evaluation with Human Performance
Evaluation models are essential for aligning generative models with human preferences. Traditional metrics like FID heusel2017gans and CLIP scores radford2021learning have been widely used huang2023t2i ; huang2024vbench ; liu2024evalcrafter . To improve evaluation accuracy, recent works fischer2020nicer ; kirstain2023pick ; liang2024rich ; wu2023humanv2 ; xu2024imagereward ; zhang2024learning have evolved from simple metrics to learning-based approaches that leverage human preference datasets to train CLIP-based models. Recently, researchers have begun exploring vision-language models (VLMs) achiam2023gpt ; wang2024qwen2 as a more powerful framework for reward modeling. These VLM-based approaches have shown success in both evaluation he2024videoscore ; wu2023humanv2 ; wu2023human and optimization lee2023aligning ; prabhudesai2024video ; wallace2024diffusion ; xu2024imagereward , utilizing methods like point-wise regression he2024videoscore ; xu2024visionreward , pair-wise comparison with Bradley-Terry loss bradley1952rank , and instruction tuning li2023generative ; lin2024criticbench ; wang2024lift to leverage reasoning capabilities for VLMs. However, VLMs need substantial data for effective training, yet expert compositional data is scarce and costly to obtain, making it challenging to train VLMs to capture sophisticated aesthetic principles using only limited expert annotations.
@@ 3.2 Automated Construction of PPC Dataset
[Intuition.] Currently, no dedicated dataset exists for PPC setting. To promote this area, we propose a novel approach for constructing PPC dataset in an automated way. The main challenge lies in collecting real-world camera movement sequences that transition from less favorable to well-composed perspective. The closest existing data comes from photographers sharing point-of-view (POV) recordings of their shooting process on streaming platforms. However, these videos are typically captured from secondary angles using GoPro, which differs from the main camera perspective. We observe that expertly composed photographs are readily available unsplash , and recent advances in single-image scene reconstruction have shown remarkable results yu2024viewcrafter . This inspired us to explore an alternative approach: generating camera movement sequences that transition from well-composed to less favorable perspective based on existing expertly composed photographs. By reversing these sequences, we can obtain the desired data.
@@
[Detail.] (1) Data Source. We select multiple professional photography datasets, including datasets used in existing composition studies such as GAIC zeng2020grid , SACD yang2023focusing , FLMS flms , and FCDB fcdb . Furthermore, to expand our data volume, we incorporated Unsplash unsplash2023 , currently the largest open-source professional photography dataset. (2) Perspective Transformation Generation. As shown in Fig. 2, we adopt a 3D reconstruction approach. Our 3D reconstruction methodology mainly builds upon the ViewCrafter yu2024viewcrafter . The inputs consist of a well-composed image and a specified camera motion trajectory. Note that this trajectory can be random (refer to [Discussion]). By following this trajectory, we can generate a video sequence transitioning from the well-composed to less favorable perspective. Then, by reversing this video sequence, we obtain our desired training data. (3) Data Filtering. Given the limited performance of current reconstruction models, the generated video data needs to filter out artifacts including distortion, fixedness, and blur effects, as depicted in Fig. 2. However, manual filtering for such a large dataset is impractical. Our tests showed that a single person can only filter about 3K videos per day, making it difficult to process large-scale samples. With the rapid advancement of vision language models (VLMs) in scene understanding and automated evaluation liu2025improving ; prabhudesai2024video ; wu2023human , we develop a perspective quality assessment (PQA) model to filter the generated data. For specific details about the PQA construction, please refer to Sec. 3.4. Our PQA evaluates generated data across three dimensions: visual quality (VQ), motion quality (MQ), and composition aesthetic (CA). These individual scores are aggregated into a final score, and samples exceeding a threshold are selected as our training data. We implement a comprehensive grading assessment that converts model-generated scores into standardized grade scores. It employs a five-tier grading scale (A to E) with corresponding numerical ranges, as shown in Tab. 1.
@@
Table 1: Grade assessment in data filtering.
A Excellent $\geq 5.0$ 95 (90-100) Superior quality
B Good 0.0 to 4.9 85 (80-89) Above-average quality
C Satisfactory -5.0 to -0.1 75 (70-79) Acceptable quality
D Marginal -15.0 to -5.1 65 (60-69) Below-average quality
E Unsatisfactory <15.0<-15.0 50 (<60<60) Inadequate quality
[Discussion.] How to handle cases where the initial perspective is less favorable? In our pipeline, we initially treat the original image as the well-composed perspective. This assumption is later refined through human preference learning during the PQA filtering (Sec. 3.4) and RLHF (Sec. 3.3) stages, acknowledging that the original perspective may not always be the aesthetically pleasing.
@@ 3.3 Photography Perspective Composition (PPC)
[Intuition.] Previous image composition works primarily focused on cropping approaches. In contrast, we propose a video-based approach: given a less favorable perspective, we generate a camera movement sequence that gradually transitions to an aesthetically enhanced perspective of the scene. This video-based approach is motivated by two key observations: First, image composition represents a partial order relationship where quality assessment often relies on comparing different views of the same scene rather than isolated evaluation. Second, the transition process naturally demonstrates compositional improvements through before-and-after comparisons, making it particularly effective for both visual demonstration and educational purposes.
[Detail.] (1) Base Pipeline. As shown in Fig. 3, our pipeline takes a less favorable perspective as input and generates a transformation video from the less favorable to aesthetically enhanced perspective. This process can be modeled as an image-to-video (I2V) task. I2V has seen remarkable progress, with both commercial solutions like OpenAI’s Sora openaisora2024 , Runway Gen-3 runway2024gen3 , and Pika 1.5 pika2024pika , and open-source models like Hunyuan kong2024hunyuanvideo , CogVideo yang2024cogvideox , and WAN wang2025wan . Our pipeline leverages these open-source models to generate perspective transformation videos. We utilize the last frame of the video as our final aesthetically enhanced perspective and design a method to guide human actions. First, we draw a guidance box (the red bbox in Fig. 3) on the enhanced perspective. Then, based on this box, along with the initial and final perspectives, we transform this box onto the original image using feature matching, creating a distorted box. As the user moves, this box gradually changes shape, approaching a rectangle when reaching the aesthetically enhanced perspective. To simplify the process and accelerate computation, we only use traditional homography transformation dubrofsky2009homography .
@@
(2) RLHF for Quality Enhancement. We observed that some generated perspective transformation videos, while deviating from the GT direction, still maintain high aesthetic quality. This suggests that strict GT adherence may not always yield the most aesthetically pleasing results. Therefore, we propose incorporating direct preference optimization (DPO) to align the model with human preferences. This approach encourages the exploration of aesthetically pleasing trajectories that may differ from GT, avoiding the limitation where GT-based optimization could discourage potentially superior compositional alternatives.
@@ 3.4 Perspective Quality Assessment (PQA) Model
[Detail.] (1) Dateset Setting. Stage ①: Unpaired Videos. This stage focuses on distinguishing video quality levels. We collected approximately 5K perspective transformation videos generated by 3D reconstruction models, with expert annotators identifying roughly 1.5K high-quality and 3.5K low-quality samples. To expand the dataset, we randomly paired each high-quality video with 10 low-quality ones, creating a 15K unpaired dataset. Stage ②: Paired Videos. This stage focuses on composition aesthetic recognition through paired comparison learning. For each initial perspective, we generate three video clips using separately trained CogVideoX 1.5, WAN 2.1, and the original GT data. These clips are paired with each other, where "paired" indicates videos sharing identical input views. Expert annotators evaluate these pairs across three dimensions: visual quality (VQ), motion quality (MQ), and composition aesthetic (CA). For each dimension, annotators choose between options (A wins/Ties/B wins). Notably, the CA metric assesses the compositional improvement throughout the video transformation rather than static frame quality.
@@
(2) Model Setting. Our model primarily follows the architecture of VideoAlign liu2025improving . We employ Qwen 2-VL as our base model, utilizing the Bradley-Terry model with ties loss (BTT) rao1967ties , which extends the traditional Bradley-Terry frameworkbradley1952rank . To better handle multi-dimensional evaluation, we separate special tokens for context-agnostic (visual quality, motion quality) and composition-aware (composition aesthetic) attributes, leveraging the causal attention mechanism to achieve effective feature decoupling. The model predicts rewards for each dimension through a shared linear projection head applied to the corresponding token representations from the final layer.
@@ 4.1 Investigation for Photography Perspective Composition (PPC)
Figure 4: Quality and human performance results for PPC.
Detailed annotation guidelines are provided in Appendix Prompt Instructions.
Metric Design. Since there is no prior work on photography composition using perspective transformation, we need to define metrics. We evaluate the generated videos from three parts: perspective accuracy, video quality, and human performance score. For video quality assessment, we primarily adopt the evaluation metrics from VBench2.0 huang2024vbench I2V benchmarks, which include I2V subject, I2V background subject consistency, background consistency, motion smoothness, dynamic degree, aesthetic quality, and image quality. For perspective accuracy evaluation, we employed both video distance and image similarity metrics. The video distance was measured using camera motion matching (CMM) and Fréchet Video Distance (FVD) unterthiner2019fvd , while image similarity was assessed using CLIP-F radford2021learning , PSNR, SSIM, and LPIPS. Notably, we modified the camera motion section of the original vbench huang2024vbench , changing it to camera motion matching. As we found that the camera motion detection model in vbench did not work well, especially for small angle changes. Instead, we now pass both the prediction and ground truth through the camera motion detection and then calculate the accuracy by matching these two detection results." As for the human performance score, its purpose is to simulate human preferences in scoring. This is conducted through PQA, which we established in Sec. 3.4, and includes three components: visual quality (VQ), motion quality (MQ), and composition aesthetic (CA).
@@
Main Results. [Quantitative Results] As shown in Tab. 2, we demonstrate the performance of three I2V models in generating PPC videos. [Qualitative Results] We demonstrate the versatility of our approach across three representative scenarios. The first scenario addresses single subjects (e.g., human figures and animals), as shown in Fig. 5, where our PPC model enhances compositional harmony by seamlessly integrating subjects with their surroundings. In the second scenario (Fig. 7), we tackle multi-subject scenes, demonstrating how PPC achieves balanced spatial arrangements to elevate overall visual aesthetics. The third scenario explores landscape photography (Fig. 8), addressing two common challenges faced by novice photographers: balance and horizontal alignment. Our model effectively optimizes these scenes, particularly enhancing symmetrical compositions. Beyond these scenarios, we discovered the applicability of PPC to UAV photography. As illustrated in Fig. 9, our model successfully identifies aesthetically enhanced views from drone-like perspectives, generating camera movements that adhere to compositional principles while maintaining aesthetic appeal.
@@
The Consistency of PPC. Fig. 6b demonstrates our model’s consistency in PPC. When presented with different less favorable views of the same scene, our model generates consistent aesthetically enhanced perspectives, maintaining coherence across different inputs.